OpenStudy (jango_in_dtown):

Complex analysis question Let D be the open unit disc. Which of the following are correct?

OpenStudy (jango_in_dtown):

i) There exists a holomorphic function f : D-> D with f(0)=0 and f'(0)=2 ii)There exists a holomorphic function f : D-> D with f(3/4)=3/4 and f'(2/3)=3/4 iii)There exists a holomorphic function f : D-> D with f(3/4)=-3/4 and f'(3/4)=-3/4 iv)There exists a holomorphic function f : D-> D with f(1/2)=-1/2 and f'(1/4)=1

OpenStudy (jango_in_dtown):

@ganeshie8

OpenStudy (jango_in_dtown):

@Loser66

OpenStudy (jango_in_dtown):

@Kainui

OpenStudy (jango_in_dtown):

@Jaynator495

OpenStudy (jango_in_dtown):

@mathstudent55

OpenStudy (jango_in_dtown):

@loser66

OpenStudy (loser66):

I saw this problem yesterday and I was so lazy to go over it. You can understand that I don't know how to do. hehehe

OpenStudy (jango_in_dtown):

I have no idea of solving this type of problem. However I read a theorem "Riemann mapping theorem", let me state it here/

OpenStudy (loser66):

which chapter are you in?

OpenStudy (jango_in_dtown):

If D is a simple connected region which is not the whole complex plane and let a belongs to D, then there exists a unique analytic function f: D-> C , with the property: i) f(a)=0 and f'(0)>0 ii) f is one-one iii) f(D)={z : |z|<1}

OpenStudy (jango_in_dtown):

But this theorem is not applicable may be

OpenStudy (loser66):

so, zero of analytic function, right?

OpenStudy (jango_in_dtown):

yeah...

OpenStudy (loser66):

I am sorry. I don't know. :)

OpenStudy (jango_in_dtown):

@ganeshie8

OpenStudy (jango_in_dtown):

@mathmate

OpenStudy (mathmate):

@jango_IN_DTOWN Sorry, not my field!

OpenStudy (jango_in_dtown):

@imqwerty

OpenStudy (holsteremission):

The details escape me at the moment (I'll need a hefty review session) but I believe this is needed here: https://en.wikipedia.org/wiki/Schwarz_lemma#Schwarz.E2.80.93Pick_theorem